DonaldTrumpofQC:

Solve this radical term:

DonaldTrumpofQC:

\[-3\sqrt{18}+3\sqrt{8}-\sqrt24\]

DonaldTrumpofQC:

I'm stuck

DonaldTrumpofQC:

i keep getting \[-9\sqrt{2}+6\sqrt{2}+-2\sqrt{6}\] , but the denominators aren't the same.

DonaldTrumpofQC:

@timmyspu

DonaldTrumpofQC:

@laylalyssa please I need help. :..(

Laylalyssa:

@donaldtrumpofqc wrote:
i keep getting \[-9\sqrt{2}+6\sqrt{2}+-2\sqrt{6}\] , but the denominators aren't the same.
this is right tho, so now u just combine like terms

DonaldTrumpofQC:

@laylalyssa wrote:
@donaldtrumpofqc wrote:
i keep getting \[-9\sqrt{2}+6\sqrt{2}+-2\sqrt{6}\] , but the denominators aren't the same.
this is right tho, so now u just combine like terms
so \[-3\sqrt{2}+-2\sqrt{6}\] , and that's it?...

Laylalyssa:

yeah good job gang

DonaldTrumpofQC:

@laylalyssa wrote:
yeah good job gang
thx